Problem epi2 SNF 2008-08-10 13:51:41: LL B door

maurice at snf.stanford.edu maurice at snf.stanford.edu
Thu Oct 2 11:17:43 PDT 2008


Door working now.  Logic could have been reset when system was powered down to pull LL pump.




More information about the epi2-pcs mailing list